OpenStudy (anonymous):

Factor this polynomial: 12x + 8

OpenStudy (anonymous):

GCF:4 \[4(\frac{12x}{4}+\frac{8}{4})=4(3x+2)\]

OpenStudy (anonymous):

I say change it into the form \[b(x +c)\] This way it is clear that the polynomial is zero when x = -c. There is little point to doing what Outkast posted.

OpenStudy (anonymous):

@Outkast3r09 can you please explain how you did this problem to me haah

OpenStudy (anonymous):

changing it completely would also yield fractions, so that'd just make it worse I used the GCF

OpenStudy (anonymous):

@LikeLightning "I say change it into the form b(x+c)" He did. And he showed HOW he did it. Your dismissal of that is unwarranted.

OpenStudy (anonymous):

The GCF (aka Greatest Common Factor) is the largest common multiples of the variables/constants that you're trying to simplify in this case, both do not have an x variable therefore x will not be in your GCF. However, This equation happens to have common factors for (12 and 8). This common factor is 4 since \[4*2=8\] and \[4*3=12\]

OpenStudy (anonymous):

so this means you can pull out a 4! But don't forget that when pulling out a factor, you're not just putting a 4 in front. That would be multiplying it by 4. Basically you're multiplying by 1 in the form of \[\frac{4}{4}=1\] \[\frac{4}{4}(12x+8)=4(\frac{12x}{4}+\frac{8}{4})\]